Be sure to answer all parts. Give your answers to the correct number of significant figures. Liquid nitrogen is obtained from liquefied air and is used industrially to prepare frozen foods. It boils at 77.36 K. (a) What is this temperature in ° C? ° C (b) What is this temperature in ° F? ° F (c) At the boiling point, the density of the liquid is 809 g/L and that of the gas is 4.566 g/L. How many liters of liquid nitrogen are produced when 358.0 L of nitrogen gas is liquefied at 77.36 K?
Answer #1

(a) 77.36 K = (77.36 - 273.15) oC = -195.79 oC

(b) F = 9C/5 + 32 = 9*(-195.79)/5 + 32 = -320.422 oF

(c) The volume of liquid nitrogen produced = 358 L * 4.566 g/L = 1634.628 g

= (1634.628 g)/(809 g/L)

= 2.02 L